OpenStudy (anonymous):

find a^1 for the geometric sequence with a^3 = 720 and a^5 = 259.2

OpenStudy (anonymous):

a^3 = a^1 * r^2 = 720 a^5 = a^1 * r^4 = 259.2 Then do simultaneous equation and solve for r then find a

OpenStudy (anonymous):

thank ya

OpenStudy (anonymous):

No Probs. Have a great day